# Non-Atomic Arbitrage in Decentralized Finance

### 15 Jan 2024 v2

@@ 5 Identifying Non-Atomic Arbitrage Trades
With our acquired understanding of non-atomic arbitrage, we now describe our procedure to identify non-atomic arbitrage trades on the Ethereum blockchain. We start by noting that identifying this type of MEV transaction is more difficult and less clear-cut than other types of MEV such as sandwich attacks, cyclic arbitrage or liquidations previously measured on Ethereum blockchain [17, 18, 19, 1, 7]. All transactions related to these types of MEV can be found on the same blockchain. Importantly, we only observe one side of the arbitrage on the Ethereum blockchain and have practically no visibility into the other side of the arbitrage. Presumably, as mentioned previously the other side of the arbitrage executes on a CEX, as these have the most significant liquidity for cryptocurrency trading. Thus, we do not have access to de-anonymized data and the detail of data available also differs highly between CEX. Note, that it is also possible that the other side of the arbitrage executes on a different blockchain, but here we are also not able to link the identity of wallets on different chains.
@@
Heuristic 3. The transaction includes a coinbase transfer to the fee recipient or its priority fee exceeds 1 GWei.
@@
Heuristic 5. The swap exchanges two established tokens, i.e., traded on CEXes (cf. Appendix E for a list).
Heuristic 1 ensures that the transaction only executes a “simple” swap on a DEX and that we do not mix it up with another type of high-value MEV transaction. Note, that we also limit the size of the transaction in units of gas to ensure that the transaction does not execute further complicated logic we have no insights into. Together, heuristics 2 and 3 allow us to identify that the transaction is valuable as it was submitted privately and tips the validator. With heuristic 4 we ensure that the transaction is the first transaction to execute in the respective direction in the pool. If this were not to be the case, the price in the pool would already have moved unfavorably for the transaction, and the sought-after arbitrage transaction might already have been closed. Finally, heuristic 5 ensures that there is likely to be another market for the trade executed, i.e., that it is possible to make the other side of the trade on a CEX for instance. Note that a transaction that violates heuristics 2, 3, 4, or 5 might still be a non-atomic arbitrage, but in that case, is likely to be a less sought-after opportunity. However, wherever possible we are conservative with what we label as non-atomic arbitrage. We also note that a transaction that does not perform any arbitrage could fulfill all our heuristics.
@@
When considering the integrated searchers from the three builders known to perform non-atomic arbitrage, we notice that our heuristics apply to nearly 90% of transactions of mantasearcher and the rsyncsearchers, but only around 60% of transactions from the beaversearchers. More than 90% transactions from mantasearcher and rsynchsearchers meet each of our heuristics. Interestingly, mantasearcher tips through coinbase transfers whereas rsyncbuild tips through priority fees. For transactions from beaversearchers, our heuristics do not apply as commonly. They are especially less likely to be private (heuristic 2), the first swap that executes in a pool (heuristic 4), and less likely to swap two established tokens (heuristic 5). When analyzing transactions by beaverbuild that do not meet the heuristic, we find that sometimes it appears that they still perform the non-atomic arbitrages, but that these are less valuable, i.e., swap less known tokens and/or only transfer smaller volumes. Thus, in these cases, it might also be less important for beaverbuild to protect their transaction, i.e., do not send the transaction privately. In other cases, it appears that beaverbuild is not performing non-atomic arbitrages, failed at performing the non-atomic arbitrage or again only performs a less valuable non-atomic arbitrage, as at least one swap that moves the price in an unfavorable direction for beaverbuilds swap precedes it. Finally, we find that for one month, the beaversearchers were not paying tips, but were instead subsidized by the beaverbuilder (cf. Appendix A.2). Thus, exclusively for beaversearchers, we do not require them to fulfill heuristic 3, if the swap executes in a block built by the beaverbuilder. In doing so, 60.4% of swaps by the beaversearchers fulfill these tailored heuristics. We note that we likely do not capture all non-atomic arbitrage transactions from beaversearchers as our heuristics are potentially too strict, but stick to our heuristic to avoid overstating the prevalence of this type of arbitrage.
@@ 6.1 Non-Atomic Arbitrage Trade Landscape
We start by providing a general overview of the non-atomic arbitrage on DEXes. On average, we identify 0.346 trades as non-atomic arbitrages per block during our data collection window that spans a little over a year. This figure is in comparison to 0.732 sandwich attacks, 0.175 cyclic arbitrage (i.e., atomic on-chain arbitrage), and 0.002 liquidations. Recall, that sandwich attacks, cyclic arbitrage, and liquidations are considered the most common and well-studied types of MEV. Thus, non-atomic arbitrage, in terms of the number of occurrences is the second most prevalent type of MEV out of these four on the Ethereum blockchain.
@@
Figure 5: Cumulative volume of non-atomic arbitrage swaps by searcher. In total, the non-atomic arbitrage volume during our data collection period is 132 billion US\$. We provide a mapping from the searcher address to the searcher name in Table VI in Appendix D.
@@
In Figure 5, we consider the cumulative volume of non-atomic arbitrage over time and highlight the eleven biggest searchers performing non-atomic arbitrage by volume. Note that we provide a mapping from searcher name to address in Appendix D. Further note that some searchers (beaversearcher1, beaverseacher2, rsyncseacher1, rsyncseacher2, rsyncseacher3, and mantaseacher) have names that suggest that they are linked to large builders in the PBS scheme (beaverbuild, rsyncbuilder and mantabuilder) [20]. We note here that this is due to our suspect that the respective builders and searchers are controlled by the same entities and we will detail our reasoning in Section 6.3.
@@
Returning to Figure 5, in total, the volume of trades we identify as non-atomic arbitrage are responsible for 132 billion US\$ on DEXes, e.g., Uniswap V2, Uniswap V3, Sushiswap, Curve, and Balancer. Startlingly, the total volume on these five DEXes in the same period of time is 460 billion US\$ [21]. Thus, the identified non-atomic arbitrage accounts for nearly 30% of the total volume of the five biggest DEXes on the Ethereum blockchain.
@@
We reiterate here that there is no clear-cut way of identifying non-atomic arbitrage, as we lack data transparency regarding the off-chain. However, the fact that merely eleven searchers are responsible for 80% of the non-atomic arbitrage volume identified in our analysis and that we identify the majority of their transaction as non-atomic arbitrage trades for all but one searcher (cf. Appendix C), leads us to believe that our heuristics are well-suited to identify non-atomic arbitrage transactions. Note that these eleven searchers can further be clustered into eight entities, three searchers associated with rsyncbuilder and the two searchers associated with beaverbuild (cf. Section 6.3 for a detailed description of the searcher/builder relationship identification). Thus, a very small number of entities is responsible for the vast majority of the non-atomic arbitrage volume on DEXes. In fact, the two biggest searchers, i.e., beaversearcher1 and jumpsearcher, account for 49.2% non-atomic arbitrage volume. We further point out that many of these large searchers (beaversearchers, jumpsearcher, rsyncsearchers, and mantasearcher) are rumored to be linked to HFT firms operating in traditional finance [4, 22], a further indicator that our heuristics capture non-atomic arbitrage as well as a sign that big players from traditional finance turning to DeFi for profits.
@@
In regards to Figure 5, we conclude that the non-atomic arbitrage volume on DEXes is immense and controlled by a few large entities. To gain a better understanding of the evolution over time, we take a more in-depth look at the share of non-atomic arbitrage volume controlled by the large searchers in Figure 6. We start by noting that on nearly 75% of days, merely two searchers are responsible for 50% of the arbitrage volume – underlying the high concentration in the non-atomic arbitrage market. Interestingly, beaverseacher1, the biggest non-atomic arbitrage searcher, is the only major searcher operating through our entire data collection period. Generally, beaversearcher1 accounts for at least 10% of the volume except for during the two time periods characterized by very high cryptocurrency price volatility in the previous, i.e., mid-November 2022 and mid-March 2023. Potentially, the non-atomic arbitrage market becomes more competitive during days characterized by exceptional events in the blockchain ecosystem, and beaverseacher1 loses some market share as a result. The second biggest searcher, jumpsearcher, on the other hand, was operating since the merge but stopped in late July 2023, while searcher1, the third biggest searcher, only started operations in mid-March 2023 and from then on was operating throughout our entire data collection window. Interestingly, the three searchers associated with rsyncbuilder all operate during non-overlapping time windows, it appears that one searcher is replacing the other – potentially upgrades to the searcher smart contract. Furthermore, there was a period of time between May and September 2023, when no searcher we could associate with rsyncbuilder was operating.
@@
TABLE III: Proportion of trades and proportion of value for each of the top searchers in pools between ETH, BTC, USDC, USDT, and DAI.
Figure 7: Distribution of the trade size of non-atomic arbitrage transactions of each of the eleven biggest searches in terms of volume. The average trade size across these eleven searchers is 73,891 US\$.
To get a better understanding of the differences in strategies utilized by the various searchers, we plot the distribution of trade sizes for each searcher as a violin plot in Figure 7. A violin plot includes a boxplot, which indicates the lower quantile, median, and upper quantile, as well as a kernel density plot to visualize the distribution of values. We observe clear differences in the trade size distribution between searchers (cf. Figure 7). For instance, beaversearcher1’s trade size distribution is quite smooth with the peak and median being around 10,000 US\$ and trade sizes ranging from less than 100 US\$ to more than 10 million US\$. Interestingly, except beaverseacher2, all other major searchers have larger median trade sizes for their non-atomic arbitrage swaps. We further note that while their median trade sizes are slightly higher than that of beaverbuild, mantaseacher and builder1searcher exhibit similar trade size distributions to beaversearcher1. A couple of searchers (e.g., jumpsearcher, searcher1, searcher2, and searcher 3) appear to be specialized in high-volume non-atomic arbitrage trades. The median trade sizes of the former three all exceed 100’000 US\$. Finally, we observe that the three rsyncbuilders tend to have several peaks in their trade size distributions, i..e, they seem to be biased towards several distinct trade sizes. Further, we observe that with each iteration of the rsyncsearcher, the median trade size declines. While we cannot be sure what the cause of this is, it could be a sign that with time rsync is turning towards smaller and likely less fought-after (as they are less valuable) non-atomic arbitrage transactions.
@@ 6.2 Drivers of Non-Atomic Trades
Looking at Figure 8, we observe that days with a high volume of non-atomic arbitrage trades tend to correspond to days with high BTC and ETH price volatility. We also observe again that the highest volume days correspond to two periods identified earlier: the FTX collapse in November 2022 and the USDC depeg in March 2023. We further note that the correlation between the daily ETH price volatility and the volume of non-atomic arbitrage is 0.725 with a p-value of $1.30\cdot 10^{-69}$ – inferring extremely high statistical significance of the correlation. For the BTC price volatility, the correlation is even higher at 0.729 with a p-value of $2.86\cdot 10^{-68}$. Thus, price volatility of ETH and BTC on CEXes drives non-atomic arbitrage on DEXes.
@@ 6.3 Integrated Seachers
To start, we analyze what proportion of the non-atomic arbitrage volume by each of the biggest eleven searchers is included in blocks built by each of the biggest nine builders (cf. Figure 10). Note that we identify the biggest searchers and builders by the non-atomic arbitrage volume they traded and included in their blocks respectively. We further provide a mapping from searcher/builder name to their address/public key in Appendix D.
@@
Similarly, 83% of the non-atomic arbitrage volume of mantasearcher is found in blocks built by the mantabuilder, while 88% of volume from the rsyncsearcher2 and an astonishing 98% of volume from the rsyncsearcher3 is included in blocks by the rsyncbuilder. Note that there is no link between rsyncsearcher1 and rsyncbuilder in Figure 10 even though our naming suggests there to be a link. We are however, able to establish a link between rsyncsearcher1 and rsyncsearcher2, as the bytecode of their contracts is identical and they are thus likely to be the same entity as established in previous work [19]. We further comment that rsynbuilder was largely not operating at the same time as rsyncsearcher1 (cf. Figure 13(b)). Finally, we find a fourth, previously not identified, builder that is likely running an integrated non-atomic arbitrage builder: builder1. To be precise, 83% of the non-atomic arbitrage volume by builder1seacher is located in blocks built by builder1seacher.

### 08 Apr 2024 v3

@@ 5 Identifying Non-Atomic Arbitrage Trades
With our acquired understanding of non-atomic arbitrage, we now describe our procedure to identify non-atomic arbitrage trades on the Ethereum blockchain. We start by noting that identifying this type of MEV transaction is more difficult and less clear-cut than other types of MEV such as sandwich attacks, cyclic arbitrage or liquidations previously measured on Ethereum blockchain [18, 19, 20, 1, 7]. All transactions related to these types of MEV can be found on the same blockchain. Importantly, we only observe one side of the arbitrage on the Ethereum blockchain and have practically no visibility into the other side of the arbitrage. Presumably, as mentioned previously the other side of the arbitrage executes on a CEX, as these have the most significant liquidity for cryptocurrency trading. Thus, we do not have access to de-anonymized data and the detail of data available also differs highly between CEX. Additionally, it is also possible that the other side of the arbitrage executes on a different blockchain, but here we are also not able to link the identity of wallets on different chains.
@@
Heuristic 3. The transaction includes a coinbase transfer to the fee recipient or its priority fee is at least 1 GWei.
@@
Heuristic 5. The swap exchanges two established tokens, i.e., traded on CEXes (cf. [17] for a list).
Heuristic 1 ensures that the transaction only executes a “simple” swap on a DEX and that we do not mix it up with another type of high-value MEV transaction. [^3] Note that we limit the size of the transaction in units of gas to ensure that the transaction does not execute further complicated logic we have no insights into. Without the gas limit we picked up transactions rebalancing liquidity positions on Uniswap V3. The limit excludes these transactions but is large enough for ordinary swaps on all five DEXes. A simple swap on in Curve V2 pools consumes the highest amount of gas for the DEXes we tracked at $\approx$350,000 gas, whereas rebalancing a liquidity position on Uniswap V3 uses $\approx$450,000 gas. Together, heuristics 2 and 3 allow us to identify that the transaction is valuable as it was submitted privately and tips the validator. [^4] For heuristic 3, we do not impose a minimum on the value of the coinbase transfer as less than 5% of transaction make such a transfer. Additionally, we set the minimum priority fee limit to 1 GWei (i.e., approximately the median priority fee of DEX swaps) to ensure that the transaction is paying a significant tip. We do not set the limit higher as beaverseachers (previously identified non-atomic arbitrage searchers) use a priority fee of 1 Gwei for more than 50% of their transactions without coinbase transfers. With heuristic 4 we ensure that the transaction is the first transaction to execute in the respective direction in the pool. If this were not to be the case, the price in the pool would already have moved unfavorably for the transaction, and the sought-after arbitrage transaction might already have been closed. Finally, heuristic 5 ensures that there is likely to be another market for the trade executed, i.e., that it is possible to make the other side of the trade on a CEX for instance. Note that a transaction that violates heuristics 2, 3, 4, or 5 might still be a non-atomic arbitrage, but in that case, is likely to be a less sought-after opportunity. However, wherever possible we are conservative with what we label as non-atomic arbitrage. We also note that a transaction that does not perform any arbitrage could fulfill all our heuristics.
@@
When considering the integrated searchers from the three builders known to perform non-atomic arbitrage, we notice that our heuristics apply to nearly 90% of transactions of mantasearcher and the rsyncsearchers, but only around 60% of transactions from the beaversearchers. Interestingly, mantasearcher tips through coinbase transfers whereas rsyncbuild tips through priority fees. Transactions from beaversearchers are especially less likely to be private (heuristic 2), the first swap that executes in a pool (heuristic 4), and less likely to swap two established tokens (heuristic 5). When analyzing transactions by beaverbuild that do not meet the heuristic, we find that sometimes it appears that they still perform the non-atomic arbitrages, but that these are less valuable, i.e., swap less known tokens and/or only transfer smaller volumes. Thus, in these cases, it might also be less important for beaverbuild to protect their transaction, i.e., do not send the transaction privately. In other cases, it appears that beaverbuild is not performing non-atomic arbitrages, failed at performing the non-atomic arbitrage or again only performs a less valuable non-atomic arbitrage, as at least one swap that moves the price in an unfavorable direction for beaverbuilds swap precedes it. Finally, we find that for one month, the beaversearchers were not paying tips, but were instead subsidized by the beaverbuilder (cf. Appendix A.2). Thus, exclusively for beaversearchers, we do not require them to fulfill heuristic 3, if the swap executes in a block built by the beaverbuilder. In doing so, 60.4% of swaps by the beaversearchers fulfill these tailored heuristics. We note that we likely do not capture all non-atomic arbitrage transactions from beaversearchers as our heuristics are potentially too strict, but stick to our heuristic to avoid overstating the prevalence of this type of arbitrage.
@@ 6.1 Non-Atomic Arbitrage Trade Landscape
We start by providing a general overview of the non-atomic arbitrage on DEXes. On average, we identify 5216 trades as non-atomic arbitrages per day during our data collection window that spans a little over a year. This figure is in comparison to 5219 sandwich attacks (front- and back-running transaction counted as one), 1245 cyclic arbitrage (i.e., atomic on-chain arbitrage), and 14 liquidations. Recall, that sandwich attacks, cyclic arbitrage, and liquidations are considered the most common and well-studied types of MEV. Thus, non-atomic arbitrage, in terms of the number of occurrences is the second most prevalent type of MEV out of these four on the Ethereum blockchain. We provide a more detailed comparison between non-atomic arbitrage and these other types of MEV in Appendix C.
@@
Figure 5: Cumulative volume of non-atomic arbitrage swaps by searcher. In total, the non-atomic arbitrage volume during our data collection period is \$132 billion. We provide a mapping from the searcher address to the searcher name as part of our data set [17].
@@
In Figure 5, we consider the cumulative volume of non-atomic arbitrage over time and highlight the eleven biggest searchers performing non-atomic arbitrage by volume. We provide a mapping from searcher name to address in our data set [17]. Further note that some searchers (beaversearcher1, beaverseacher2, rsyncseacher1, rsyncseacher2, rsyncseacher3, and mantaseacher) have names that suggest that they are linked to large builders in the PBS scheme (beaverbuild, rsyncbuilder and mantabuilder) [21]. This naming is due to our suspicion that the respective builders and searchers are controlled by the same entities and we will detail our reasoning in Section 6.3.
@@
Returning to Figure 5, in total, the volume of trades we identify as non-atomic arbitrage are responsible for \$132 billion on DEXes, e.g., Uniswap V2, Uniswap V3, Sushiswap, Curve, and Balancer. Startlingly, the total volume on these five DEXes in the same period of time is \$460 billion [22]. Thus, the identified non-atomic arbitrage accounts for nearly 30% of the total volume of the five biggest DEXes on the Ethereum blockchain.
@@
We reiterate here that there is no clear-cut way of identifying non-atomic arbitrage, as we lack data transparency regarding the off-chain. However, the fact that merely eleven searchers are responsible for 80% of the non-atomic arbitrage volume identified in our analysis and that we identify the majority of their transaction as non-atomic arbitrage trades for all but one searcher (cf. Appendix D), leads us to believe that our heuristics are well-suited to identify non-atomic arbitrage transactions. Additionally, these eleven searchers can further be clustered into eight entities, three searchers are associated with rsyncbuilder and two searchers are associated with beaverbuild (cf. Section 6.3 for a detailed description of the searcher/builder relationship identification). Thus, a very small number of entities is responsible for the vast majority of the non-atomic arbitrage volume on DEXes. In fact, the two biggest searchers, i.e., beaversearcher1 and jumpsearcher, account for 49.2% non-atomic arbitrage volume. We further point out that many of these large searchers (beaversearchers, jumpsearcher, rsyncsearchers, and mantasearcher) are rumored to be linked to HFT firms operating in traditional finance [4, 23], a further indicator that our heuristics capture non-atomic arbitrage as well as a sign that big players from traditional finance turning to DeFi for profits.
@@
In regards to Figure 5, we conclude that the non-atomic arbitrage volume on DEXes is immense and controlled by a few large entities. To gain a better understanding of the evolution over time, we take a more in-depth look at the share of non-atomic arbitrage volume controlled by the large searchers in Figure 6. We start by noting that on nearly 75% of days, merely two searchers are responsible for at least 50% of the arbitrage volume – underlying the high concentration in the non-atomic arbitrage market. Interestingly, beaverseacher1, the biggest non-atomic arbitrage searcher, is the only major searcher operating through our entire data collection period. Generally, beaversearcher1 accounts for at least 10% of the volume except for during the two time periods characterized by very high cryptocurrency price volatility in the previous, i.e., mid-November 2022 and mid-March 2023. Potentially, the non-atomic arbitrage market becomes more competitive during days characterized by exceptional events in the blockchain ecosystem, and beaverseacher1 loses some market share as a result. The second biggest searcher, jumpsearcher, on the other hand, was operating since the merge but stopped in late July 2023, while searcher1, the third biggest searcher, only started operations in mid-March 2023 and from then on was operating throughout our entire data collection window. Interestingly, the three searchers associated with rsyncbuilder all operate during non-overlapping time windows, it appears that one searcher is replacing the other – potentially upgrades to the searcher smart contract. Furthermore, there was a period of time between May and September 2023, when no searcher we could associate with rsyncbuilder was operating.
@@
TABLE III: Proportion of trades and value for each of the top searchers in pools between ETH, BTC, USDC, USDT, and DAI.
Figure 7: Distribution of the trade size of non-atomic arbitrage transactions of each of the eleven biggest searchers in terms of volume. The average trade size across these eleven searchers is \$73,891.
To get a better understanding of the differences in strategies utilized by the various searchers, we plot the distribution of trade sizes for each searcher as a violin plot in Figure 7. A violin plot includes a boxplot, which indicates the lower quantile, median, and upper quantile, as well as a kernel density plot to visualize the distribution of values. We observe clear differences in the trade size distribution between searchers (cf. Figure 7). For instance, beaversearcher1’s trade size distribution is quite smooth with the peak and median being around \$10,000 and trade sizes ranging from less than \$100 to more than \$10 million. Interestingly, except beaverseacher2, all other major searchers have larger median trade sizes for their non-atomic arbitrage swaps. We further note that while their median trade sizes are slightly higher than that of beaverbuild, mantaseacher and builder1searcher exhibit similar trade size distributions to beaversearcher1. A couple of searchers (e.g., jumpsearcher, searcher1, searcher2, and searcher 3) appear to be specialized in high-volume non-atomic arbitrage trades. The median trade sizes of the former three all exceed \$100,000. Finally, we observe that the three rsyncbuilders tend to have several peaks in their trade size distributions, i..e, they seem to be biased towards several distinct trade sizes. Further, we observe that with each iteration of the rsyncsearcher, the median trade size declines. While we cannot be sure what the cause of this is, it could be a sign that with time rsyncsearcher is turning towards smaller and likely less fought-after (as they are less valuable) non-atomic arbitrage transactions.
@@ 6.2 Drivers of Non-Atomic Trades
Looking at Figure 8, we observe that days with a high volume of non-atomic arbitrage trades tend to correspond to days with high BTC and ETH price volatility. Additionally, the highest volume days correspond to two periods identified earlier: the FTX collapse in November 2022 and the USDC depeg in March 2023. We further note that the correlation between the daily ETH price volatility and the volume of non-atomic arbitrage is 0.725 with a p-value of $1.30\cdot 10^{-69}$ – inferring extremely high statistical significance of the correlation. For the BTC price volatility, the correlation is even higher at 0.729 with a p-value of $2.86\cdot 10^{-68}$.
@@
Note that the strong correlation between price volatility and non-atomic arbitrage volume does not necessarily imply causation. To further explore relationship, we calculate the correlation between volume (excluding non-atomic arbitrage) and non-atomic arbitrage volume. The correlation is 0.655, and thus, lower than the correlation between non-atomic arbitrage volume and price volatility of 0.725 and 0.729 for ETH and BTC respectively. The stronger correlation between non-atomic arbitrage volume and price volatility, underlines the relationship between the two. Additionally, we note that the correlation is lower despite our heuristics being conservative and likely undercounting the prevalence of non-atomic arbitrage, i.e., some remaining volume is probably non-atomic arbitrage and thereby correlated. Thus, price volatility of ETH and BTC on CEXes likely drives non-atomic arbitrage on DEXes.
@@ 6.3 Integrated Seachers
To start, we analyze what proportion of the non-atomic arbitrage volume by each of the biggest eleven searchers is included in blocks built by each of the biggest nine builders (cf. Figure 10). Note that we identify the biggest searchers and builders by the non-atomic arbitrage volume they traded and included in their blocks respectively. We provide a mapping from searcher/builder name to their address/public key in our data set [17].
@@
Similarly, 83% of the non-atomic arbitrage volume of mantasearcher is found in blocks built by the mantabuilder, while 88% of volume from the rsyncsearcher2 and an astonishing 98% of volume from the rsyncsearcher3 is included in blocks by the rsyncbuilder. Note that there is no link between rsyncsearcher1 and rsyncbuilder in Figure 10 even though our naming suggests there to be a link. However, we are able to establish a link between rsyncsearcher1 and rsyncsearcher2, as the bytecode of their contracts is identical and they are thus likely to be the same entity as established in previous work [20]. We further comment that rsynbuilder was largely not operating at the same time as rsyncsearcher1 (cf. Figure 13(b)). Finally, we find a fourth, previously not identified, builder that is likely running an integrated non-atomic arbitrage builder: builder1. To be precise, 83% of the non-atomic arbitrage volume by builder1seacher is located in blocks built by builder1seacher.